OpenStudy (jordanloveangel):

1.How many solutions does the equation have? 4(2x+3)=2(3x-4) A. one solution B. infinite solutions C. no solution

OpenStudy (jordanloveangel):

2. (1 pt) How many solutions does the equation have? A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

number 1 is A

OpenStudy (yanasidlinskiy):

Use Distributive Property first. You get \(\huge{8x+12=6x-8}\) Do you know what is the next step?

OpenStudy (yanasidlinskiy):

@queenoftherings Direct Answers are AGAINST the CoC. Do NOT give out direct answers instead, help the person out:) Thank you.

OpenStudy (anonymous):

oh alright

OpenStudy (yanasidlinskiy):

Just a heads up. I don't any of the users to get suspended just because of that:) @jordanloveangel do you know what is the next step?

OpenStudy (jordanloveangel):

no

OpenStudy (yanasidlinskiy):

\(\huge{8x+12=6x-8}\) \(\huge{-6x -6x}\) \(\huge{2x+12=-8}\) Add 12 to both sides. What do you get? 2x= ?

OpenStudy (jordanloveangel):

-10 ?

OpenStudy (yanasidlinskiy):

My mistake, it's supposed to -12 from both sides.

OpenStudy (yanasidlinskiy):

Signs are the same so you add:) What do you get?

OpenStudy (jordanloveangel):

but do u still get -10

OpenStudy (yanasidlinskiy):

No. You would add -12 + -8

OpenStudy (jordanloveangel):

-20?@YanaSidlinskiy

OpenStudy (jordanloveangel):

@The_Jokers_wife

OpenStudy (anonymous):

what @jordanloveangel

OpenStudy (yanasidlinskiy):

Yep! So now you have: \(\huge{2x=-20}\) Divide both sides by 2. What do you get? If it's no solution then it means that Both sides are unequal. So you can rule out that one:)

OpenStudy (jordanloveangel):

dam can u help

OpenStudy (jordanloveangel):

for get it joker

OpenStudy (jordanloveangel):

A right

OpenStudy (yanasidlinskiy):

Nope. Try again?

OpenStudy (jordanloveangel):

b then right

OpenStudy (yanasidlinskiy):

Yes:)

OpenStudy (anonymous):

wait wait if you plug in 1 for example your not gonna get them to equal each other @YanaSidlinskiy

OpenStudy (yanasidlinskiy):

nope.